Estimate the following for a nitrogen molecule. The diameter of a nitrogen molecule is approximately 0.300 nm.

Estimate the mean free path of a nitrogenmolecule in air at the top of Mt. Everest (altitude = 8.80 km, P ≈ 50.0 kPa, and T ≈ 248 K). Value of Boltzmann's constant is 1.381×10−23 J/K

Answer must be in nm. please show process. Thank you!
Answer #1

Mean free path path 1. RTKBT n= n= EndNAP rendre a=8-314 X 248 v2 = x(0:3)x108x6.023x1o yooxió 9 = 5 38 nm 5 38 nm